Analysis of the data from 25 mothers indicates that an infants birth weight (g), y. can be estimated by a mothers weight (kg), x, using the regression equation y-31x+1501. 13) Interpret the slope of the regression equation. 13)
Answer #1

Though there could be a number of other factors that might influence birth weight, this simplistic regression model (y = 31x + 1501) emphasizes that an infant's birth weight seems to be directly proportional to the weight of the mother.

The slope 31 signifies that the infant's birth weight (dependent variable) seems to increase by 31 gms for per kg increase in mother's weight (independent variable).

E.g., if mother's weight goes up from 60->61 kg, infant birth weight goes up from y = 31*60+1501 = 3361 gms, to y = 31*61+1501 = 3392 gms, an increase of 31 gms in agreement with the slope of the regression model.
Answer #2

The answer is 3671

source: COURSE HERO
answered by: DEAL